### v1

@@ B.4 Full Pass-Through Elasticity
This section extends the analysis to incorporate certain indirect effects by capturing how a shock to exporter $i$ influences prices and quantities in other relationships, and how those changes feed back into the bilateral price $p_{ij}$. While we continue to abstract from full general equilibrium forces, this exercise clarifies how network spillovers may cause reduced-form pass-through estimates to deviate from structural ones, as emphasized by Berger et al. (2022).
We generalize the cost shock to an exporter-level shock, denoted $\vartheta_{i}$, and re-derive the relevant elasticities to allow for cross-relationship spillovers. First, the impact on the importer’s buyer share becomes:
@@
$\frac{d\ln c_{i}}{d\ln\vartheta_{i}}$ $=\frac{1-\theta}{\theta}\left(-\varepsilon_{ij}x_{ij}\frac{d\ln p_{ij}}{d\ln\vartheta_{i}}-\sum_{z\in\mathcal{Z}_{i},\,z\neq j}x_{iz}\varepsilon_{iz}\frac{d\ln p_{iz}}{d\ln\vartheta_{i}}\right).$
@@ C.1 Related-Party Trade Measured via Ownership Linkages
Using these same rules, we replicate the MID structure for foreign firms in the ORBIS database. We then match each MID in the customs data to candidate firms in ORBIS based on the reconstructed name segment. We assess the quality of each potential match using two dimensions: location and product alignment. A location score is computed based on the match between city names in the MID and in ORBIS. A product match score is computed by comparing the NAICS6 industry code listed in ORBIS to the HS6 product code recorded in the customs data, using the concordance developed by Pierce and Schott (2009). We retain only those matches where both location and product scores exceed 90%. In addition, we drop from the matched dataset any ORBIS firm with fewer than five transactions to filter out spurious exporters and potential noise.
@@
Taken together, these steps yield a linked dataset that offers a more transparent and conservative definition of related-party trade, based on majority ownership (at least 50%). In contrast to the standard related-party trade (RPT) flag in customs data, which applies a lower threshold of 6% ownership for imports, this approach reduces false positives and more precisely captures transactions where ownership ties are likely to influence pricing. By combining MID-based matching with firm-level ownership structures, the final dataset is well-suited for analyzing pricing behavior in cross-border transactions.
@@ C.2 Data Cleaning, Sample Construction and Summary Statistics
We construct the analysis sample in several steps to align the data with the model’s structure and requirements.
We begin by removing observations that are incomplete or inconsistent with the modeling framework. Specifically, we drop transactions with missing or zero values for import value or quantity, invalid exporter identifiers (e.g., strings with fewer than three characters or beginning with a number), or U.S. importers that cannot be linked to the Longitudinal Business Database (LBD). We also exclude transactions associated with special provisions or temporary classifications (HS codes 98–99).
Next, we restrict our attention to trade in intermediate and capital goods by removing HS10 products, which are classified as consumption goods under the BEC system.
Although the customs data are reported at the transaction level, we aggregate them to the annual level for each buyer–supplier–product triplet. We choose annual aggregation because few relationships appear in adjacent months or in the same month across years. We then retain only those triplets that are active in at least two consecutive years and where the supplier transacts the same HS10 product with more than one U.S. buyer. This ensures a panel structure with repeated observations, supporting the identification of the model’s parameters.
We also exclude related-party trade, which is less likely to reflect decentralized bargaining and more likely to involve internal pricing practices such as transfer pricing. [^35] Bernard et al. (2006) shows that related-party prices differ systematically from arm’s-length transactions, including lower average prices and different pass-through behavior. In our baseline, a buyer–supplier pair is flagged as related if ORBIS identifies a shared corporate parent. To preserve coverage, we retain all observations not flagged as related by either ORBIS or the LFTTD. For robustness, we also consider two alternative definitions: one based solely on the LFTTD flag, and another combining it with ORBIS data to identify the U.S. importer as a multinational. See Appendix C.1 for further details.
@@
To address outliers, we follow Heise (2024) and apply two filters. First, we drop observations with extreme price levels, defined as log unit values below the 1st or above the 99th percentile of the HS10 product–country distribution. Second, we trim extreme price changes by excluding year-on-year log price differences smaller than $-4$ or greater than $+4$. These outliers are removed from estimation but retained when computing tenure and relationship length. We also exclude all HS10 products under HS chapter 27 (energy-related goods), which lie outside the scope of the model.
@@
After applying these restrictions, which exclude outliers, energy products, and related-party trade, and focusing on repeated, arm’s-length relationships, the sample retains over 20% of U.S. imports by value and nearly 15% of buyer–supplier–product triplets.
Finally, we restrict attention to suppliers that sell the same HS10 product to more than one U.S. buyer. This condition is essential for the identification of the model’s firm-level parameters and ensures that the empirical setting aligns with the model’s structure oulined in Setion 4.
@@
The final rows show the estimation samples used in the analysis. For 2001–2016 (Panel A), the data include approximately \$880 billion in import value, 480,000 buyer–supplier pairs, and 630,000 buyer–supplier–HS10 triplets. For 2017–2018 (Panel B), the final sample includes \$160 billion in imports, 190,000 pairs, and 250,000 triplets. [^36] Panel A omits intermediate sample steps and shows only the full and final samples for 2001–2016, as these are not reported in the analysis nor disclosed under Census data policies. The same applies to Table C.2. These samples correspond exactly to the data used in estimation and post-estimation analysis.
Table C.2 reports a similar analysis using a broader sample that includes capital, intermediate, and consumption goods (“BEC – All (Consec.)”), rather than limiting to consumption goods only. Analogous to Table 1, Table C.3 reports the summary statistics of this broader sample.
@@
Panel A: Characteristics of Trade Relationships
@@
Panel B: Prices
@@
Notes: This table reports summary statistics for a sample that covers all BEC product categories except energy goods, and excludes statistical outliers and related-party trade. It further restricts to suppliers that trade with at least two different U.S. buyers in consecutive years. This corresponds to the cumulative sample underlying the “+ Supplier Multi-Buyer” row in Panel B of Table C.2. Columns report the mean, standard deviation, and selected quantiles (25th percentile, median, and 75th percentile) for each variable. Prices in Panel B are log unit values (FOB value over quantity), with variants including charges or duties. $s_{ijh}$ denotes exporter $i$’s share in buyer $j$’s imports of product $h$; $x_{ijh}$ denotes buyer $j$’s share in exporter $i$’s U.S. exports of the same product. Relationship length and tenure are in years; concentration is measured at the HS10–year level. Counts of buyers, suppliers, and origin countries are per product per firm. Statistics are based on confidential LFTTD data and rounded to four significant digits per U.S. Census Bureau Disclosure Guidelines. Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
@@ D.1 Decomposition of Price Dispersion
estimated over the period 2001-2016. We consider three alternative prices: prices that exclude both duties and charges ($\ln p_{ijht}$), prices that include charges but exclude duties ($\ln p^{c}_{ijht}$), and prices that include duties but exclude charges ($\ln p^{\text{duty}}_{ijht}$).
@@
$FE_{j}$ 0.231 0.233 0.233
Match residual 0.768 0.765 0.765
Notes: The columns correspond to alternative price definitions: $\ln p_{ijht}$ excludes both duties and charges; $\ln p^{c}_{ijht}$ includes charges but excludes duties; $\ln p^{\text{duty}}_{ijht}$ includes duties but excludes charges. The estimation sample includes importer–exporter–product matches observed in two consecutive calendar years, and applies the following restrictions: (i) excludes transactions involving consumption goods (based on the BEC classification), energy-sector products, statistical outliers, and related-party trade; and (ii) retains only suppliers that trade with at least two distinct U.S. buyers in consecutive years. This corresponds to the cumulative sample underlying the “+ Supplier Multi-Buyer” row in Panel B of Table C.1. The control vector $\boldsymbol{X}_{ijht}$ includes the log of transaction value, the log of relationship longevity (years since the exporter first supplied the buyer with the given HS10 product), and the log of the relative number of partners (the supplier’s number of HS10-level buyers divided by the buyer’s number of HS10-level suppliers). The sample includes 1,2000,000 importer–exporter–product–year observations, which have been rounded to four significant digits per U.S. Census Bureau disclosure guidelines. $R^{2}=0.956$. All coefficients in a regression model are significantly different from zero at the 1% significance level. Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).

### v2

@@ B.4 Full Pass-Through Elasticity
This section extends the analysis to incorporate certain indirect effects by capturing how a shock to exporter $i$ influences prices and quantities in other relationships, and how those changes feed back into the bilateral price $p_{ij}$. While we continue to abstract from full general equilibrium forces, this exercise clarifies how network spillovers may cause reduced-form pass-through estimates to deviate from structural ones, as emphasized by berger2022labor.
We consider an exporter-level price shifter, denoted $\vartheta_{i}$, and re-derive the relevant elasticities to allow for cross-relationship spillovers. First, the impact on the importer’s buyer share becomes:
@@
$\frac{d\ln MC_{i}}{d\ln\vartheta_{i}}$ $=\frac{1-\theta}{\theta}\left(-\varepsilon_{ij}x_{ij}\frac{d\ln p_{ij}}{d\ln\vartheta_{i}}-\sum_{z\in\mathcal{Z}_{i},\,z\neq j}x_{iz}\varepsilon_{iz}\frac{d\ln p_{iz}}{d\ln\vartheta_{i}}\right).$
@@
B.5 Buyer Power, Decreasing Returns, and Pass-Through
We compare a perfectly competitive benchmark with decreasing returns to scale to our bilateral bargaining framework in order to clarify the source of pass-through heterogeneity. Consider one supplier $i$ and two buyers $j\in\{1,2\}$. The supplier’s marginal cost is
$MC_{i}=k_{i}q_{i}^{\frac{1-\theta}{\theta}},$
which implies a marginal cost elasticity
$mc_{q}\equiv\frac{d\ln MC_{i}}{d\ln q_{i}}=\frac{1-\theta}{\theta}>0.$
Total output is $q_{i}=q_{i1}+q_{i2}$ and buyer shares are defined as $x_{ij}\equiv q_{ij}/q_{i}$. For simplicity, both buyers have constant demand elasticity $\sigma\equiv-d\ln q_{ij}/d\ln p_{ij}$.
Under perfect competition and in the absence of price discrimination, prices are uniform and equal to marginal cost, $p_{ij}=p_{i}=MC_{i}$. A tariff shock $T$ enters multiplicatively into marginal cost, so that $d\ln p_{i}=d\ln MC_{i}+d\ln T$. Using $d\ln MC_{i}=mc_{q}\,d\ln q_{i}$ together with $d\ln q_{i}=-\sigma\,d\ln p_{i}$, we obtain
$d\ln p_{i}=-mc_{q}\sigma\,d\ln p_{i}+d\ln T,$
which implies
$\Phi^{\text{PC}}\equiv\frac{d\ln p_{i}}{d\ln T}=\frac{1}{1+mc_{q}\sigma}=\frac{1}{1+\frac{1-\theta}{\theta}\sigma}.$
Pass-through is therefore incomplete due to decreasing returns and identical across buyers.
In the bargaining model, prices are set bilaterally according to $p_{ij}=\mu_{ij}MC_{i}$. Pass-through becomes relationship-specific:
$\Phi^{\text{B}}\equiv\frac{d\ln p_{ij}}{d\ln T}=\frac{1}{1+\Gamma_{ij}+\Lambda_{ij}},$
where $\Gamma_{ij}$ captures markup adjustments and
$\Lambda_{ij}\equiv mc_{q}\,x_{ij}\,\sigma$
captures the elasticity of marginal cost induced by buyer $j$’s share of output.
Quantitatively, under our estimate of $\theta$, the markup channel is dominated by the cost channel (see Figure 2), so that $\Gamma_{ij}\approx 0$. Pass-through is then well approximated by
$\Phi^{\text{B}}\approx\frac{1}{1+\Lambda_{ij}}=\frac{1}{1+\frac{1-\theta}{\theta}\sigma x_{ij}}.$
The two environments share the same underlying mechanism: incomplete pass-through is governed by the slope of marginal cost, $mc_{q}$. The difference lies in the unit of adjustment. Under perfect competition, the relevant elasticity is aggregate and pass-through is uniform. Under bargaining, the effective elasticity is share-weighted, generating heterogeneity across relationships. Large buyers internalize a steeper effective supply curve and exhibit lower pass-through, whereas small buyers approximate the constant-returns benchmark. Bargaining thus provides a micro-foundation for decreasing-returns-type pass-through while delivering the relationship-level heterogeneity observed in the data.
@@ C.1 Related-Party Trade Measured via Ownership Linkages
Using these same rules, we replicate the MID structure for foreign firms in the ORBIS database. We then match each MID in the customs data to candidate firms in ORBIS based on the reconstructed name segment. We assess the quality of each potential match using two dimensions: location and product alignment. A location score is computed based on the match between city names in the MID and in ORBIS. A product match score is constructed at the MID–firm level by comparing the foreign firm’s NAICS6 industry code in ORBIS to the distribution of HS6 products shipped under that MID in the customs data, using the concordance developed by Pierce2009. We retain only those matches where both the location and product match scores exceed 90%. In addition, we drop from the matched LFTTD–ORBIS dataset any ORBIS-linked firm that is associated with fewer than five customs transactions. This filter is based on the number of shipment records matched to each ORBIS identifier in the merged data and is intended to eliminate weak or potentially noisy MID–firm matches.
@@
Taken together, these steps yield a linked dataset that offers a more transparent and conservative definition of related-party trade, based on majority ownership (at least 50%). In contrast to the standard related-party trade (RPT) flag in customs data, which applies a lower threshold of 5% ownership for imports, this approach reduces false positives and more precisely captures transactions where ownership ties are likely to influence pricing. By combining MID-based matching with firm-level ownership structures, the final dataset is well-suited for analyzing pricing behavior in cross-border transactions.
@@ C.2 Data Cleaning, Sample Construction and Summary Statistics
We construct the analysis sample in several steps to align the data with the model and the identification strategy. As in the main text, it is useful to distinguish between restrictions that are essential for identification and those that primarily ensure consistency with the model’s environment.
Preliminary restrictions.
We begin by aggregating transaction-level customs records to the annual importer–exporter–product (HS10) level. Because relatively few relationships appear in adjacent months or in the same month across years, annual aggregation provides a natural unit of observation for relationship-level price changes.
@@
To ensure reliable price measurement, we apply several data-quality filters. We drop transactions with missing or zero import value or quantity, invalid exporter identifiers, or U.S. importers that cannot be linked to the Longitudinal Business Database (LBD). We exclude HS codes 98–99 (special provisions), remove observations with log unit values below the 1st or above the 99th percentile of the HS10–country distribution, and trim extreme year-on-year log price changes outside the interval [4,4][-4,4]. [^33] These filters are applied to the estimation sample but retained when computing tenure and relationship length. We also exclude energy goods (HS chapter 27), which fall outside the scope of the model.
Identification-driven restrictions.
First, we require importer–exporter–product triplets to be observed in at least two consecutive calendar years. This panel structure is necessary to compute year-on-year price changes and to estimate pass-through.
Second, our identification strategy relies on cross-sectional price variation across buyers sourcing the same product from the same supplier. We therefore restrict attention to supplier–product pairs that transact with at least two distinct U.S. buyers in consecutive years. This condition is essential for identifying the bargaining and returns-to-scale parameters, as outlined in Section 4.
Model-alignment restrictions.
The remaining restrictions ensure that the empirical setting matches the model’s focus on decentralized bargaining over intermediate inputs. We exclude products classified as consumption goods under the Broad Economic Categories (BEC) system, restricting attention to capital and intermediate inputs.
We also exclude related-party trade, as such transactions may reflect intra-firm transfer pricing rather than arm’s-length bargaining.343434bernard2006transfer document systematic differences between related-party and arm’s-length prices. In the baseline, a buyer–supplier pair is classified as related if ORBIS identifies a common corporate parent. To preserve coverage, we retain all observations not flagged as related by either ORBIS or the LFTTD. For robustness, we consider alternative definitions based on the LFTTD related-party flag and on multinational status (see Appendix C.1).
@@
After applying these restrictions, the final sample isolates repeated, arm’s-length buyer–supplier relationships in intermediate goods with sufficient cross-sectional variation for identification; it retains over 20% of U.S. imports by value and nearly 15% of buyer–supplier–product triplets.
@@
The final rows report the estimation samples used in the analysis. For 2001–2016 (Panel A), the sample includes approximately \$880 billion in import value, 480,000 buyer–supplier pairs, and 630,000 buyer–supplier–HS10 triplets. For 2017–2018 (Panel B), the corresponding sample comprises \$160 billion in imports, 190,000 pairs, and 250,000 triplets. [^35] Panel A omits intermediate sample steps and reports only the full and final samples for 2001–2016, as intermediate counts are neither used in the analysis nor releasable under Census disclosure policies. The same applies to Table C.3. These samples correspond exactly to those used in the structural estimation and post-estimation analysis. Table C.2 reports summary statistics for the main estimation sample, analogous to Table 1 in the main text.
Tables C.3 and C.4 present analogous results for a broader sample that includes capital, intermediate, and consumption goods (“BEC – All (Consec.)”), rather than restricting to non-consumption goods.
@@
Table C.2: Summary Statistics for Main Estimation Sample (2001–2018)
Variable Mean Std. Dev. P25 Median P75
Panel A: Trade Relationships
   $s_{ijh}$: Supplier share 0.32 0.35 0.03 0.15 0.57
   $x_{ijh}$: Buyer share 0.25 0.29 0.02 0.10 0.40
   Relationship length (product $h$) 4.00 2.80 2.50 3.50 5.50
   Relationship length (all products) 4.80 3.30 2.50 4.50 6.50
   # Transactions (product $h$) 120 1100 6.50 16 50
   # Transactions (all products) 360 3000 11 36 140
   # Products per pair 3.80 7.30 1.50 2.50 4.50
   Multi-HS10 dummy 0.59 0.49 0.00 1.00 1.00
   # Suppliers per buyer (HS10) 1.80 3.20 1.50 2.50 5.50
   Buyer tenure (all products) 9.90 5.00 6.50 10.00 14.00
   Buyer tenure (product $h$) 6.90 4.40 3.50 6.50 10.00
   # Buyers per supplier (HS10) 3.20 3.90 2.50 3.50 7.50
   Supplier tenure (all products) 8.00 4.60 4.50 8.50 12.00
   Supplier tenure (product $h$) 6.40 4.00 3.50 6.50 9.50
   Corr. between $s_{ijh}$ and $x_{ijh}$ 0.041
Panel B: Prices (log unit values)
   $\log p$ (pre-duty, excl. charges) 3.40 2.80 1.30 3.00 5.40
   $\log p$ (pre-duty) 3.50 2.80 1.40 3.10 5.40
   $\log p^{\text{duty}}$ (post-duty) 3.50 2.80 1.40 3.10 5.40
Notes: This table extends Table 1 with additional variables for the main estimation sample at the importer-supplier-product level. Full details on sample construction are provided in Appendix C.2. Panel A reports trade relationship characteristics. $s_{ijh}$ is exporter $i$’s share in buyer $j$’s imports of product $h$; $x_{ijh}$ is buyer $j$’s share in exporter $i$’s U.S. exports of product $h$. Relationship length (product $h$) is the number of years between the supplier-buyer pair’s first transaction in product $h$ and year $t$; Relationship length (all products) is defined analogously, using the pair’s first transaction in any product. # Transactions (product $h$) is the cumulative number of transactions between the pair in product $h$, from their first transaction through year $t$; # Transactions (all products) is defined analogously, summing across all products. # Products per pair is the number of distinct HS10 products traded by the pair in year $t$, and Multi-HS10 dummy equals one when this number exceeds one. # Suppliers per buyer (HS10) is the number of distinct exporters from which buyer $j$ sources product $h$ in year $t$; # Buyers per supplier (HS10) is defined analogously, counting distinct buyers per exporter. Buyer tenure (product $h$) is the number of years between the buyer’s first recorded transaction in product $h$ (with any supplier) and year $t$; Buyer tenure (all products) is defined analogously, using the buyer’s first recorded transaction in any product. Supplier tenure is defined analogously for the exporter. Corr. between $s_{ijh}$ and $x_{ijh}$ is the sample correlation between the two shares. Panel B reports log unit values (value divided by quantity). $\log p$ (pre-duty, excl. charges) corresponds to the FOB (Free on Board) unit value, excluding shipping and insurance charges; $\log p$ (pre-duty) includes these charges to obtain the CIF (Cost, Insurance, and Freight) unit value; and $\log p^{\text{duty}}$ (post-duty) further incorporates import duties. Statistics are based on confidential LFTTD data and rounded according to U.S. Census Bureau disclosure guidelines. Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
@@
Panel A: Trade Relationships
@@
Panel B: Prices (log unit values)
   $\log p$ (pre-duty, excl. charges) 3.30 2.50 1.40 3.00 5.00
@@
Notes: This table reports summary statistics for the pooled sample, defined analogously to the main estimation sample (Table 1) but without restricting on BEC end-use category, at the importer-supplier-product level. Full details on sample construction are provided in Appendix C.2. Panel A reports trade relationship characteristics. $s_{ijh}$ is exporter $i$’s share in buyer $j$’s imports of product $h$; $x_{ijh}$ is buyer $j$’s share in exporter $i$’s U.S. exports of product $h$. Relationship length (product $h$) is the number of years between the supplier-buyer pair’s first transaction in product $h$ and year $t$; Relationship length (all products) is defined analogously, using the pair’s first transaction in any product. # Transactions (product $h$) is the cumulative number of transactions between the pair in product $h$, from their first transaction through year $t$; # Transactions (all products) is defined analogously, summing across all products. # Products per pair is the number of distinct HS10 products traded by the pair in year $t$, and Multi-HS10 dummy equals one when this number exceeds one. # Suppliers per buyer (HS10) is the number of distinct exporters from which buyer $j$ sources product $h$ in year $t$; # Buyers per supplier (HS10) is defined analogously, counting distinct buyers per exporter. Buyer tenure (product $h$) is the number of years between the buyer’s first recorded transaction in product $h$ (with any supplier) and year $t$; Buyer tenure (all products) is defined analogously, using the buyer’s first recorded transaction in any product. Supplier tenure is defined analogously for the exporter. Corr. between $s_{ijh}$ and $x_{ijh}$ is the sample correlation between the two shares. Panel B reports log unit values (value divided by quantity). $\log p$ (pre-duty, excl. charges) corresponds to the FOB (Free on Board) unit value, excluding shipping and insurance charges; $\log p$ (pre-duty) includes these charges to obtain the CIF (Cost, Insurance, and Freight) unit value; and $\log p^{\text{duty}}$ (post-duty) further incorporates import duties. Statistics are based on confidential LFTTD data and rounded according to U.S. Census Bureau disclosure guidelines. Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
@@ D.1 Decomposition of Price Dispersion
estimated over the period 2001-2016. We consider three alternative prices: prices that exclude duties ($\ln p_{ijht}$), prices that exclude duties and charges ($\ln p^{c}_{ijht}$), and prices that include duties and charges ($\ln p^{\text{duty}}_{ijht}$).
@@
$FE_{j}$ 0.233 0.231 0.233
Match residual 0.765 0.768 0.765
Notes: The columns correspond to alternative price definitions: $\ln p_{ijht}$ excludes duties; $\ln p^{c}_{ijht}$ exclude both duties and charges; $\ln p^{\text{duty}}_{ijht}$ includes duties. The estimation sample includes importer–exporter–product matches observed in two consecutive calendar years, and applies the following restrictions: (i) excludes transactions involving consumption goods (based on the BEC classification), energy-sector products, statistical outliers, and related-party trade; and (ii) retains only suppliers that trade with at least two distinct U.S. buyers in consecutive years. This corresponds to the cumulative sample underlying the “+ Supplier Multi-Buyer” row in Panel B of Table C.1. The control vector $\boldsymbol{X}_{ijht}$ includes the log of transaction value, the log of relationship longevity (years since the exporter first supplied the buyer with the given HS10 product), and the log of the relative number of partners (the supplier’s number of HS10-level buyers divided by the buyer’s number of HS10-level suppliers). The sample includes 1,200,000 importer–exporter–product–year observations, which have been rounded to four significant digits per U.S. Census Bureau disclosure guidelines. $R^{2}=0.956$. All coefficients in a regression model are significantly different from zero at the 1% significance level. Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).